From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Jan 16 2003 - 08:12:00 EST
------------------------------------------------------------ Date: Thu Jan 16 08:11:59 2003, Shift: 00:00-08:00 Supervisor: AM, with: CR Log book#13 pages 67 - 68 Shift Summary: ------------------------------------------------------------ 1. No beam until 2 am. After we had some beam but with lot of background. Around 5.30 the beam was cogged and backg rates become resonable. 2. Run 7410 - not registrered into DB. Finished as we got IOServer crash. (~150k trig6, 15k trig3) Satarted run 7411 (T3 is on). 3. HV MF for TPCs went down. We still kept anodes off during high background beam. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Thu Jan 16 2003 - 08:13:47 EST